The scale of a model is 0.5 in: 6 ft. Find the length of the model for the actual length of 204 ft.
Answer by jim_thompson5910(35256)

(0.5 in)/(6 ft) = (x in)/(204 ft)

0.5/6 = x/204

0.5*204 = 6x

102 = 6x

6x = 102

x = 102/6

x = 17

So the length of the model is 17 inches.